OpenStudy (anonymous):

can someone help me plz.... I need to solve this: x+2y=8 x=4-2y

OpenStudy (anonymous):

1) ask yourself what the slopes of both equations are. Then ponder on the answer

OpenStudy (anonymous):

lol, I just don't understand how to put it together... I think 2y-x=8. see, this is what confuses me

OpenStudy (anonymous):

Well ... the answer for this problem comes from identifying the slope for the two equations of lines that you are given. The solution is where these two line cross. trust me. Please just ask yourself whether the two slopes of the two different lines are different.

OpenStudy (anonymous):

hmmmm, not understanding you, can you show me an example

OpenStudy (anonymous):

Ok EQ 1) x + 2y =8 the slope can be done by converting to slope-int form 2y = 8-x y = 4-0.5x So the slope is -0.5 EQ 2) x = 4-2y x-4 = -2y -0.5x +2 = y Slope = -0.5 So this question is basically asking you when two lines with the same slope intersect. When is that?